# Unifying the Perspectives of NLP and Software Engineering: A Survey on Language Models for Code

### v3

@@ Appendix A Benchmarks for Downstrem Tasks
Task Date Benchmark Source Size Language Code Retrieval 2018-03 StaQC Yao et al. (2018) 268K Python, SQL 2018-05 DeepCS Gu et al. (2018) 16M Java 2018-05 CoNaLa Yin et al. (2018) 600K/2.9K Python 2019-08 unnamed Li et al. (2019b) 287 Java 2019-09 CodeSearchNet Husain et al. (2019) 2.3M/99 Go, JS, Python, PHP, Java, Ruby 2020-02 CosBench Yan et al. (2020) 52 Java 2020-08 SO-DS Heyman & Cutsem (2020) 2.2K Python 2020-10 FB-Java Ling et al. (2021) 249K Java 2021-02 AdvTest Lu et al. (2021) 280K Python 2021-02 WebQueryTest Lu et al. (2021) 1K Python 2021-05 CoSQA Huang et al. (2021) 21K Python Code Reasoning 2020-09 MMLU Hendrycks et al. (2021b) 15908 2021-09 CodeQA Liu & Wan (2021) 120K/70K Java/Python 2022-10 CS1QA Lee et al. (2022) 9237 2023-05 C-Eval Huang et al. (2023b) 13948 2023-06 CMMLU Li et al. (2023a) 11528 2023-09 CodeApex Fu et al. (2023) 250 Type Inference 2019-12 TypeWriter OSS Pradel et al. (2020) 208K Python 2020-04 Typilus Allamanis et al. (2020) 252K Python 2020-04 LambdaNet Wei et al. (2020) 300 TypeScript 2021-04 ManyTypes4Py Mir et al. (2021) 869K Python 2022-10 ManyTypes4TypeScript Jesse & Devanbu (2022) 9.1M TypeScript 2023-02 TypeWeaver Yee & Guha (2023) 513 TypeScript 2023-03 BetterTypes4Py Wei et al. (2023) 608K Python 2023-03 InferTypes4Py Wei et al. (2023) 4.6K Python 2023-05 OpenTau Cassano et al. (2023b) 744 TypeScript Clone Detection / Code Search 2014-09 BigCloneBench Svajlenko et al. (2014) 6M Java 2014-09 POJ-104 Mou et al. (2016) 52K C, C++ 2019-05 unnamed Perez & Chiba (2019) 21K/24K Java/Python 2019-11 CLCDSA Nafi et al. (2019) 78K Java, C#, Python
@@
Task Date Benchmark Source Size Language Unit Test / Assertion Generation 2014-12 SF110 Fraser & Arcuri (2014) 24K Java 2020-09 Method2Test Tufano et al. (2021a) 781K Java 2021-08 ConTest Villmow et al. (2021) 365K Java Log Parsing 2018-11 LogHub (2018) Zhu et al. (2019); He et al. (2020) 379M 2023-08 LogHub (2023) Jiang et al. (2023c) 50.4M Repository-Level Coding 2023-03 RepoEval Zhang et al. (2023b) 1600/1600/373 Python 2023-06 RepoBench Liu et al. (2023i) 890K/9M/43K Python, Java 2023-06 PragmaticCode Agrawal et al. (2023) 880 Java 2023-06 Stack-Repo Shrivastava et al. (2023a) 816K Java 2023-09 CodePlan Bairi et al. (2023) 645/21 C#/Python 2023-10 SWE-Bench Jimenez et al. (2023) 2294 Python 2023-10 CrossCodeEval Ding et al. (2023) 9928 Python, Java, TypeScript, C#

### v5

@@ Appendix A Benchmarks for Downstrem Tasks
Task Date Benchmark Source Size Language Code Retrieval 2018-03 StaQC Yao et al. (2018) 268K Python, SQL 2018-05 DeepCS Gu et al. (2018) 16M Java 2018-05 CoNaLa Yin et al. (2018) ∗600K/2.9K Python 2019-08 unnamed Li et al. (2019b) 287 Java 2019-09 CodeSearchNet Husain et al. (2019) ∗2.3M/99 Go, JS, Python, PHP, Java, Ruby 2020-02 CosBench Yan et al. (2020) 52 Java 2020-08 SO-DS Heyman & Cutsem (2020) 2.2K Python 2020-10 FB-Java Ling et al. (2021) 249K Java 2021-02 AdvTest Lu et al. (2021) 280K Python 2021-02 WebQueryTest Lu et al. (2021) 1K Python 2021-05 CoSQA Huang et al. (2021) 21K Python 2024-03 ProCQA Li et al. (2024b) 5.2M C, C++, Java, Python, Ruby, Lisp, JS, C#, Go, Rust, PHP Code Reasoning 2020-09 MMLU Hendrycks et al. (2021b) †15908 2021-09 CodeQA Liu & Wan (2021) 120K/70K Java/Python 2022-10 CS1QA Lee et al. (2022) 9237 2023-05 C-Eval Huang et al. (2023b) †13948 2023-06 CMMLU Li et al. (2023b) †11528 2023-09 CodeApex Fu et al. (2023) 250 Type Inference 2019-12 TypeWriter OSS Pradel et al. (2020) 208K Python 2020-04 Typilus Allamanis et al. (2020) 252K Python 2020-04 LambdaNet Wei et al. (2020) ‡300 TypeScript 2021-04 ManyTypes4Py Mir et al. (2021) 869K Python 2022-10 ManyTypes4TypeScript Jesse & Devanbu (2022) 9.1M TypeScript 2023-02 TypeWeaver Yee & Guha (2023) ‡513 TypeScript 2023-03 BetterTypes4Py Wei et al. (2023) 608K Python 2023-03 InferTypes4Py Wei et al. (2023) 4.6K Python 2023-05 OpenTau Cassano et al. (2023c) ‡744 TypeScript Clone Detection / Code Search 2014-09 BigCloneBench Svajlenko et al. (2014) 6M Java 2014-09 POJ-104 Mou et al. (2016) 52K C, C++ 2019-05 unnamed Perez & Chiba (2019) ⋄21K/24K Java/Python 2019-11 CLCDSA Nafi et al. (2019) 78K Java, C#, Python
@@
Task Date Benchmark Source Size Language Unit Test / Assertion Generation 2014-12 SF110 Fraser & Arcuri (2014) 24K Java 2020-09 Method2Test Tufano et al. (2021a) 781K Java 2021-08 ConTest Villmow et al. (2021) 365K Java Log Parsing 2018-11 LogHub (2018) Zhu et al. (2019); He et al. (2020) 379M 2023-08 LogHub (2023) Jiang et al. (2023e) ∗50.4M Repository-Level Coding 2023-03 RepoEval Zhang et al. (2023b) †1600/1600/373 Python 2023-06 RepoBench Liu et al. (2023j) ‡890K/9M/43K Python, Java 2023-06 PragmaticCode Agrawal et al. (2023) ⋆880 Java 2023-06 Stack-Repo Shrivastava et al. (2023a) 816K Java 2023-09 CodePlan Bairi et al. (2023) ⋄645/21 ⋄C#/Python 2023-10 SWE-Bench Jimenez et al. (2023) 2294 Python 2023-10 CrossCodeEval Ding et al. (2023) 9928 Python, Java, TypeScript, C# 2024-03 EvoCodeBench Li et al. (2024a) 275 Python